# AI’s Social Forcefield: Reshaping Distributed Cognition in Human–AI Teams

### v2

@@ Abstract.
AI is not only a neutral tool in team settings; it actively reshapes the social and cognitive fabric of collaboration. We advance a unified framework of alignment in distributed cognition in human–AI teams—a process through which linguistic, cognitive, and social coordination emerge as human and AI agents co-construct a shared representational space. Across two studies, we show that exposure to AI-generated language shapes not only how people speak, but also how they think, what they attend to, and how they relate to each other. Together, these findings reveal how AI participation reorganizes the distributed cognitive architecture of teams: AI systems function as implicit “social forcefields.” Together, our findings highlight the double-edged impact of AI: the same mechanisms that enable efficient collaboration can also erode epistemic diversity and undermine natural alignment processes. We argue for rethinking AI in teams as a socially influential actor and call for new design paradigms that foreground transparency, controllability, and group-level dynamics to foster responsible, productive human-AI collaboration.
Human-AI teaming, collective attention, linguistics, entrainment, shared mental models, lexical alignment
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Some; 2025; Placeisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
AI is increasingly pervasive in modern workplaces (Brynjolfsson et al., 2025; Agarwal et al., 2023) and projected to have wide ranging impact on the future of work (BCG, 2024; AI Index, 2025). Yet many are increasingly concerned about its influence (Kennedy et al., 2025; Chayka, 2025). AI use is not isolated from other goings on as workers are embedded in teams and engaged in social interactions with human coworkers. The key thesis of this paper is that AI systems do not just passively assist human teams; they actively reshape the social and cognitive fabric of teamwork. Our work shifts the view of AI from a passive tool or assistant to an actor that actively influences cognitive alignment in human teams. We develop and test an unified framework of cognitive alignment in human–AI teams that links micro-level cognitive coupling (word reuse & shared phrasing) to higher-level collective cognition (shared focus of attention and shared understanding) and social cohesion & group identity. The framework specifies how AI-induced linguistic adaptation shapes shared attention and cognitive alignment, producing what we term AI’s social forcefield. Even when the AI is no longer present, linguistic traces continue to structure how humans coordinate with each other, shaping the persistent alignment of distributed cognition in human–AI teams.
This work is a detailed exploration how human-human communication is influenced by AI. It moves the conversation from “How can AI help humans?” to “How does AI rewire human collectives, and how do we design for that responsibly?” Our investigation considers cognitive alignment as an overarching, generative framework (Figure 1). We consider how lower-level influences cascade upward, via cumulative stabilization of alignment to shape distributed cognition in teams (Ocasio, 2011; Woolley et al., 2023; Riedl et al., 2021). Shared language reflects low-level cognitive coupling, enabling individuals to coordinate their representations in real time through mutual adaptation in speech (Clark, 1996; Sievers et al., 2024; Tausczik and Pennebaker, 2010). Linguistic entrainment in turn creates the representational scaffolding for shared attention and alignment of mental models (Clark, 1996; Kommol et al., 2023; Fusaroli et al., 2012; Hollan et al., 2000). That is, by naming things the same way, groups can sustain joint focus, and start seeing the task and each other through the same conceptual lens (Tomasello, 2009; Shteynberg, 2015; DeChurch and Mesmer-Magnus, 2010; Tausczik and Pennebaker, 2010). Shared mental models and attention in turn can amplify memories, goals, emotions, and behavioral strategies (Shteynberg, 2015; Clark and Brennan, 1991; Woolley et al., 2023), allowing groups to experience a sense of cognitive fluency and affective coherence (Hutchins, 1995); the feeling that “we understand each other” thus signals group identity and social cohesion (Tausczik and Pennebaker, 2010). The AI’s linguistic contributions penetrate upward through the layers of this theoretical framework—from phrasing, to shared focus of attention and shared mental models, to social cohesion and team identity—affecting the social and cognitive fabric of groups. Even when absent, the AI remains cognitively “in the loop.”
@@
While recent work has made significant advances in understanding human-AI interaction (Rezwana and Maher, 2023; Riedl and Weidmann, 2025; Inkpen et al., 2023; Xiao et al., 2020; Bansal et al., 2019a; Branigan et al., 2011; Cowan et al., 2015; Wudarczyk et al., 2021; Zhang et al., 2023), these studies predominantly focus on dyadic relationships (one-user-one-AI) and immediate task performance (e.g., Zając et al., 2025; Vasconcelos et al., 2025; Arzberger et al., 2024; Ma et al., 2025). These studies often overlook the complexities of multi-member teams, and AI’s multifaceted effect on cognitive alignment across levels and beyond the direct interaction with AI. Prior work that is explicitly focused on teams, typically explores how teams use AI—or how AI can assist in teamwork—but often reported mixed results and high variability (e.g., Autor, 2014; Fügener et al., 2021; Dell’Acqua et al., 2023). We address these gaps by showing that AI creates a “social forcefield” that persists in human-human interaction, affecting cognitive alignment across multiple levels. Work often examines isolated AI effects (Wang et al., 2022) rather than interconnected cognitive alignment mechanisms in team settings (Cuevas et al., 2007). Even studies examining subtle AI influences, such as the placebo (Kosch et al., 2023) or ghostwriter effect (Draxler et al., 2024), investigate isolated effects during active AI engagement (Chaves et al., 2022), while we show how AI language spills over into subsequent human-human conversations and scaffolds cognitive alignment across multiple levels. As AI becomes an integral part of teamwork, addressing these challenges is critical to unlocking its potential to strengthen human-AI partnerships and guard against unintended consequences like homogenization. To address these gaps, we answer the following research questions:
@@ 1. RQ1
    Does experimentally manipulating exposure to AI language causally alter team communication (language, attention, coordination) during collaboration, and does this influence spill over into subsequent human-only interaction?
@@ 2. RQ2
    How robust are these effects across AI attributes (helpfulness and voice anthropomorphism), across interaction modalities (voice assistant and text-based chat), and across phases with and without active AI participation?
@@ 3. RQ3
    Which social‑cognitive appraisals (e.g., trust, perceived intelligence, in‑group membership) predict the magnitude of these effects?
We provide evidence from two randomized controlled studies to show that exposure to AI-generated language shapes not only how people speak, but also how they think, what they attend to, and how they relate to each other. In Study 1, 20 teams of 3-4 individuals perform a 40-minute group task. Each team was randomly assigned a voice-only AI assistant that provided helpful or unhelpful information, and spoke in a human or robotic voice. Our analysis then explores (a) the AI’s effect on teams alignment on shared language, (b) teams shared focus of attention and mental models, and (c) group identity and cohesion. In Study 2, 97 individuals used ChatGPT 4o to answer customer service complaints. During this interaction, we exposed them to subtle wording differences by instructing the chatbot (via a randomly assigned system prompt) to suggest customer service responses that emphasized either warmth and empathy or formal adherence to company guidelines. Users then reflect on their approach to the task in a team-focused face-to-face debriefing interview. We then analyze overlap between subjects’ spoken words and the AI words they were exposed to by ChatGPT during the task. This design directly tests whether subtle variations in an AI assistant’s task framing and language leaves a linguistic and cognitive “fingerprint” that persists in human-human interaction beyond the AI’s direct involvement. Across both studies, a variety of data sources, and analysis approaches we find consistent evidence that AI shapes the cognitive alignment of human-human interaction, from the language people use when communicating with other humans, to their shared attention and mental models, and social cohesion and team identity.
In summary, we make the following contributions to our understanding of how AI influences cognitive alignment across multiple layers.
@@
    We show that AI significantly influences linguistic alignment, shared attention and mental models, and group identity and social cohesion.. This happens even though the AI assistant has limited control over team interactions (Study 1) or even when it is no longer present (Study 2), holds across different types of AI (voice assistant vs. text-based chat), quality levels (high, low, or naturalistic), or when subjects report not trusting it.
@@
    We reconceptualize AI as an implicit “social forcefield”: AI systems embedded in human teams inevitably shape human-human communication patterns, collective cognition, and shared identity—whether designers intend this or not. Unlike most studies that focus on one-on-one interactions (either human-to-human or human-to-AI), we examine how AI influences the communication and thinking of human groups beyond the direct interaction with the AI.
@@
    We highlight the double-edged impact of AI: It can help align language and distributed cognition, but also shapes collective cognition when not intended to—undermining natural alignment processes—and introduces homogeneity that reduces cognitive diversity.
@@
    We call for a new paradigm: AI will inevitably reshape human–human interaction. While its effects can be bounded through transparency and controllability, they cannot be fully predefined. We argue that effective human–AI teaming requires attention not only to functional performance, but also to the social and cognitive processes through which AI systems can enhance—rather than disrupt or homogenize—collective intelligence.
2. Background
Individual minds become coupled through communication and subtle linguistic cues can redirect what groups attend to and how they relate to each other. Our hypothesis is that AI assistants might not just be tools that provide information or automate tasks; they also affect the social and cognitive fabric of teamwork. Introducing AI in settings of human-human collaboration, might shift the way people talk (through linguistic entrainment) which in turn shifts what people pay attention and how they relate to each other. We conceptualize cognitive alignment as a generative framework that links micro-level linguistic coordination to macro-level collective outcomes (Figure 1). At the micro level, linguistic entrainment—individuals reusing words and phrasing introduced by the AI—creates a shared representational substrate. This alignment in language enables meso-level convergence in attention and shared mental models, as teams begin to focus on the same issues and structure the task through similar conceptual lenses. Over time, such cognitive coherence gives rise to macro-level social integration, expressed through in-group identification (e.g., inclusive pronouns like “we”). In this view, cognitive alignment is not a collection of parallel effects but a cascade of interdependent processes through which AI-introduced language shapes how people think, what they attend to, and how they relate to one another. This framework extends distributed cognition from being computer mediated and shaped by digital objects (Hollan et al., 2000) into the social domain of human–AI teams.
Figure 1. AI exerts a “social forcefield,” affecting the cognitive alignment in human groups through a multi-level model, from Micro-level coupling of linguistic entrainment, through meso-level alignment of attention and shared mental models, to macro-level alignment through relational and social integration.:
@@ 2.1. Shared Language
On the lowest level of our framework sits shared language. Shared language reflects the micro-level cognitive coupling that allows individuals to align their representations as they adapt to each other’s speech in real time. Lexical alignment occurs when conversational partners converge on similar words or phrases during interaction (Wachsmuth et al., 2013). Over time, these repetitions form “conceptual pacts,” which are agreements on how to refer to things within a group (Brennan and Clark, 1996). Alignment on shared language reduce cognitive effort; without shared language communication would be more difficult and increase processing demands (Metzing and Brennan, 2003). Shared language therefore helps groups coordinate attention, conserve mental resources, and collaborate more efficiently (Kalociński et al., 2018).
2.2. Shared Attention and Shared Mental Models
@@
Shared attention reflects the meso-level convergence of focus, through which individuals align what aspects of the task are salient and meaningful, sustaining a common frame of reference for joint action. Language carries shared representations and structuring how people think and act together. Shared language alignment is thus the linguistic foundation of collective attention which shows priorities and intentions (Tausczik and Pennebaker, 2010). Human attention is the ability to actively process specific information while tuning out other, less relevant details (Bundesen, 1990). Attention is limited in both capacity and duration (Duncan, 2016), making its effective coordination essential for high performance (Duncan, 2013). In group contexts, individuals align their attention with others, including what they perceive, remember, and prioritize (Kampis and Southgate, 2020). Groups achieve high collective attention when members co-attend to the same object or event, creating a shared psychological experience that shapes motivation, judgment, and behavior (Shteynberg, 2015; Woolley et al., 2023). Collective attention thus enables teams to coordinate activities and sustain joint action (Mayo and Woolley, 2021).
Collective attention is often studied through multiple lenses: the temporal focus of groups on shared information (e.g., social media activity) (Weng et al., 2012), synchrony in behaviors like facial expression or vocal rhythm (Woolley et al., 2023), and, crucially, the alignment of language (Sievers et al., 2024). Language can play a central role because it links cognitive processes to shared meaning, aligning beliefs and neural activity across team members (Nagai et al., 2003; Kaplan and Hafner, 2006). By shaping how problems are described and discussed, collective attention also influences abstract intentions, plans, and strategies (Hawkins et al., 2023), which in turn can affect creativity and performance (Laureiro-Martinez et al., 2023; Bergenholtz et al., 2023).
@@
Together with shared attention, shared mental models are the collective cognitive structures that teams develop to interpret and predict their environment (DeChurch and Mesmer-Magnus, 2010). They capture a group’s shared understanding of “what is happening, what is likely to happen next, and why it is happening” (Mohammed et al., 2010, p. 879). Such models improve motivation, coordination, and performance, particularly when tasks require mutual dependence on diverse skills and expertise (Lewis, 2003). Their development relies on the temporal alignment of attention and the establishment of shared language, which together provide the building blocks for joint understanding (Tollefsen et al., 2013; DeChurch and Mesmer-Magnus, 2010; Mohammed et al., 2010). That is, the shared lexicon on the micro-layer scaffolds shared understanding of the broader context of work through the development of shared mental models and joint focus of attention.

### v3

# Cognitive Spillover in Human-AI Teams

@@ Abstract.
AI is not only a neutral tool in team settings; it influence the social and cognitive fabric of collaboration. Across two randomized experiments, we demonstrate that AI exposure produces causal spillover into human-human interaction—affecting shared language, collective attention, shared mental models, and social cohesion. These spillover effects occur robustly across settings, modalities, tasks, and AI qualities, suggesting that mere exposure to AI drives the influence. AI functions as an implicit “social forcefield,” influencing not only how people speak, but also how they think, what they attend to, and how they relate to each other. We argue for shifting the design paradigm from optimizing “AI as a tool” to understanding AI as a socially influential actor whose effects extend beyond the human-AI interface.
Human-AI teaming, team cognition, collective attention, linguistics, entrainment, shared mental models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Some; 2025; Placeisbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Psychologyccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing HCI theory, concepts and modelsccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Multi-agent systems
@@ 1. Introduction
Artificial Intelligence (AI) is increasingly pervasive in modern workplaces (Brynjolfsson et al., 2025; Agarwal et al., 2023) and projected to have wide ranging impact on the future of work (BCG, 2024; AI Index, 2025). Yet there is also growing unease about indirect consequences (Kennedy et al., 2025; Chayka, 2025) and heightened regulatory interest in systemic AI risk (Bengio et al., 2024). In particular, AI use does not occur in isolation as workers are embedded in teams where they are engaged in social interactions with human coworkers. The key thesis of this paper is that AI systems do not just passively assist users; they influence the social and cognitive fabric of teamwork beyond their direct interactions with the AI. By shifting the analytical lens from “AI as tool” to “AI as social forcefield,” we show how AI influence can spill over to human-human interaction beyond AI’s immediate functional role.
We draw on prior literature on distributed cognition (Hutchins, 1995; Hollan et al., 2000) to articulate how AI influences cognitive alignment in human–AI teams (Pickering and Garrod, 2004b; Garrod and Pickering, 2009). A core insight from this literature is that alignment operates simultaneously across multiple coupled channels: linguistic, attentional, representational, and affective (Garrod and Pickering, 2009; Clark, 1996). Shared language enables groups to coordinate how they understand situations in real time (Clark, 1996; Sievers et al., 2024; Tausczik and Pennebaker, 2010); collective attention orients groups toward common objects and topics (Kommol et al., 2025; Woolley et al., 2023; Riedl and Woolley, 2017; Ocasio, 1997); shared mental models align interpretive frameworks and behavioral expectations (Fusaroli et al., 2012; Hollan et al., 2000; DeChurch and Mesmer-Magnus, 2010); and the affective sense that “we understand each other,” shapes group identity and social cohesion (Shteynberg, 2015; Tomasello, 2009; Hutchins, 1995). While alignment at one level can promote alignment at others, these channels operate in parallel rather than in strict causal sequence (Pickering and Garrod, 2004b). These constructs span the depth of collaborative dynamics, from automatic (linguistic) to deliberative (cognitive) to affective (social), capturing how the introduction of AI can affect coupled alignment processes within human teams.
A major challenge in measuring AI’s influence in team settings is distinguishing between direct human-AI alignment and spillover into subsequent human-human interaction, while controlling for the natural alignment resulting from shared task history alone. It is well known that people align with their current conversation partner’s language (e.g., word choice, syntax, style) (Wachsmuth et al., 2013; Brennan and Clark, 1996; Garrod and Anderson, 1987), even when that partner is an AI (Branigan et al., 2011; Hu et al., 2014). It is also known that interaction dynamics themselves naturally produce shared language even without explicit interpersonal influence (Garrod and Anderson, 1987; Brennan and Clark, 1996; Clark and Wilkes-Gibbs, 1986). To claim that AI actively shapes team cognition, one must rigorously disentangle causal AI influence from the natural baseline of task-driven vocabulary alignment and look beyond the (dyadic) influence during AI use itself.
@@
Current HCI literature has not yet resolved this challenge. While recent work has made significant advances in understanding human-AI interaction (Rezwana and Maher, 2023; Riedl and Weidmann, 2025; Inkpen et al., 2023; Xiao et al., 2020; Bansal et al., 2019a; Branigan et al., 2011; Cowan et al., 2015; Wudarczyk et al., 2021), prior work focuses largely on effects within the human-AI dyad during active use (e.g., Zając et al., 2025; Vasconcelos et al., 2025; Ma et al., 2025), without testing whether such effects produce spillover into human-human dynamics. Team-focused research reports mixed results—AI is sometimes helpful, sometimes harmful, often unpredictably variable (e.g., Autor, 2014; Fügener et al., 2021; Dell’Acqua et al., 2026)—in part because it typically examines isolated outcomes like performance (Wang et al., 2022) rather than the multi-channel alignment mechanisms that coordinate teamwork (Schelble et al., 2022). And studies of subtle AI influences—such as performance boosts from perceived AI support (Kosch et al., 2023) or perceived ownership in AI-assisted writing (Draxler et al., 2024)—focus on effects during active AI engagement (Chaves et al., 2022), leaving open whether AI shapes social interaction among team members beyond their direct interaction with the AI. We address these gaps by testing for AI spillover across multiple parallel channels of socio-cognitive alignment—linguistic, attentional, cognitive, and relational—and tracing effects beyond active AI engagement into subsequent human-human interaction.
In this paper, we provide evidence of causal spillover in two randomized controlled studies while carefully controlling for shared-task baselines. Study 1 focuses specifically on isolating causal spillover of AI language to subsequent human-human interaction. Participants used ChatGPT 4o to respond to customer service complaints. By manipulating the AI’s system prompt (empathic vs. formal), we exposed users to systematic linguistic variations while holding the underlying task and interaction dynamic constant. We then analyzed users’ subsequent language use during a face-to-face debriefing interview with a human partner. This design directly tests whether subtle differences in the AI assistant’s task framing and language leave a linguistic and cognitive “fingerprint” that persists in subsequent human–human interaction beyond the shared task vocabulary. Study 2 expands this to a dynamic, ecologically complex team setting. Teams of 3–4 individuals performed a complex problem-solving task with a voice-based AI assistant. We manipulated AI attributes (helpfulness and voice anthropomorphism) and measured the effects on shared language, collective attention, shared mental models, and social cohesion.
This research answers three high-level questions:
@@ 1. RQ1
    Causal Spillover: Does exposure to AI causally affect socio-cognitive alignment of human teams—linguistic, cognitive, and social—beyond the direct interaction with the AI and beyond natural alignment resulting from shared task alone?
@@ 2. RQ2
    Robustness: Are these effects robust across different AI modalities (text vs. voice), attributes (anthropomorphism and helpfulness), and phases of activity (AI-present vs. AI-absent)?
@@ 3. RQ3
    Conscious vs. Implicit Influence: To what extent is this socio-cognitive alignment dependent on positive user appraisal (e.g., trust, perceived intelligence, perceived team membership), or does AI affect team dynamics through an implicit, automatic process regardless of user sentiment?
In summary, we make the following contributions to the understanding of how AI influences socio-cognitive alignment in human-AI teams across multiple parallel channels.
@@
    Empirical Evidence of Spillover: We provide causal evidence that AI influence spills over into human-human interaction. By rigorously controlling for shared task vocabulary, we demonstrate that this effect is not merely dyadic entrainment, but a detectable “residue” in human-human communication.
@@
    AI Influence on Multiple Channels: We show that AI-induced spillover is not confined to surface-level linguistic mimicry but manifests across multiple parallel channels of socio-cognitive alignment—from automatic linguistic coordination to deliberative shared mental models to affective group cohesion. This breadth is nontrivial: spillover could plausibly have been absent, weak, or limited to a single channel (e.g., Pickering and Garrod, 2004a; Danescu-Niculescu-Mizil et al., 2012; Glikson and Woolley, 2020; Danescu-Niculescu-Mizil et al., 2012; Gray and Wegner, 2012; Reeves and Nass, 1996).
@@
    Generalizability Across Contexts: By triangulating evidence across two heterogeneous studies (individual vs. team, text vs. voice, customer service vs. problem-solving task, high- vs. low-capacity agents), we show that these alignment effects are robust. Notably, they are detectable even when the AI is no longer present, has limited capacity, or when users report low trust, suggesting an implicit cognitive effect distinct from conscious reliance. The convergence of effects across different contexts provides stronger evidence of generalizable predictive validity than any single test (Campbell and Fiske, 1959; Cook et al., 2002; Lawlor et al., 2016).
@@
    Reconceptualization of AI in Teams: We argue for shifting the design paradigm from optimizing “AI as a tool” to understanding “AI as a social forcefield.” The spillover effects we document emerge readily and affect different parallel channels of alignment, suggesting that AI’s influence on team cognition is broader and less contained than current design frameworks assume.
2. Conceptual Framework for Socio-Cognitive Alignment in Human–AI Teams
Effective collaboration in teams—whether human-only, human-AI, or AI-AI—requires that members represent the task and each other in mutually compatible ways (Hutchins, 1995; Riedl and Weidmann, 2025; Riedl, 2026). This representational alignment enables shared reference (“what are we talking about?”), coordinated attention (“what matters right now?”), prediction of others’ behavior (“what will they do next?”), and effective coordination (“can I rely on them to interpret this the same way?”). Thus, alignment is not a single construct but a multi-channel process, observable at different levels of interaction (Hutchins, 1995; Clark, 1996; Resnick et al., 1991). We theorize that AI assistants—and especially AI language models—are not passive tools; instead, they mediate alignment by introducing linguistic and conceptual anchors that can persist beyond the direct (dyadic) interaction with the AI itself and influence the cognitive and social fabric of teamwork on different levels. Thus, our work extends insights on how digital artifacts shape coordination (Hollan et al., 2000; Hutchins, 1995) to the social dynamics of human–AI teams. [^1] When we speak of human-AI teams we mean multiple humans not a single human-AI dyad.
Grounded in prior work on distributed and team cognition (Hutchins, 1995; Clark, 1996; Resnick et al., 1991; Cooke et al., 2000), we adopt a literature-based framework that points to alignment processes happening on multiple levels in parallel (Figure 1). We then operationalize each level using established measures and test whether AI affects outcomes on that level. This literature-based framework serves as an analytical lens to structure our inquiry and inform our measures.
Figure 1. Conceptual framework of AI spillover on different levels of cognitive and social alignment. (Left) AI exerts a “social forcefield,” with its influence spilling over to human-human interactions. (Right) Three parallel channels: linguistic, cognitive, and social.
Conceptual framework of AI spillover on different levels of cognitive and social alignment.
@@ 2.1. Shared Language
Shared language reflects cognitive coupling that allows individuals to align their representations as they adapt to each other’s speech in real time. Lexical alignment occurs when conversational partners converge on similar words or phrases during interaction (Wachsmuth et al., 2013). Over time, these repetitions form “conceptual pacts,” which are agreements on how to refer to things within a group (Brennan and Clark, 1996). Alignment on shared language reduce cognitive effort; without shared language, communication would be more difficult and increase processing demands (Metzing and Brennan, 2003). Shared language therefore helps groups coordinate attention, conserve mental resources, and collaborate more efficiently (Kalociński et al., 2018; Hutchins, 1995; Ocasio, 1997).
Causally attributing spillover to AI influence requires paying careful attention to what the baseline should be (Pickering and Garrod, 2004a; Clark and Wilkes-Gibbs, 1986). Convergence on shared task vocabulary naturally arises from task structure (Garrod and Anderson, 1987), common ground (Clark and Brennan, 1991), and coordination history (Brennan and Clark, 1996). A “no influence” baseline would therefore be too lenient as it risks attributing naturally occurring alignment to AI-induced spillover. To demonstrate causal AI spillover, we need to establish effects above and beyond this baseline level of alignment (from task structure, common ground, coordination history, etc). Hence, our analyses operationalize AI spillover as the reuse of AI-specific lexical items in subsequent human-human interaction above baseline alignment on shared task vocabulary (see Methods for details). Whereas traditional forced-naming tasks may overestimate effects by capturing direct lexical substitution and anchoring (Brandstetter et al., 2017), our approach measures whether AI exposure shapes the language people naturally use, even when no explicit labeling decision is required. For example, in our second study, teams are not prompted to talk about the object mentioned by the AI but may choose to do so at their own volition.
2.2. Collective Attention and Shared Mental Models
@@
Human attention is the ability to actively process specific information while tuning out other, less relevant details (Bundesen, 1990). Attention is limited in both capacity and duration (Duncan, 2016), making its effective coordination essential for high performance (Duncan, 2013). In group contexts, individuals align their attention with others, including what they perceive, remember, and prioritize (Kampis and Southgate, 2020). Groups achieve high collective attention when members co-attend to the same object or event, creating a shared psychological experience that shapes motivation, judgment, and behavior (Shteynberg, 2015; Woolley et al., 2023). Collective attention thus enables teams to coordinate activities and sustain joint action (Mayo and Woolley, 2021). Prior work establishes that shared language plays a central role in shaping collective attention because it links cognitive processes to shared meaning, aligning beliefs and neural activity across team members (Nagai et al., 2003; Kaplan and Hafner, 2006). By shaping how tasks are described and discussed, language influences collective attention through its alignment of intentions, plans, and strategies (Hawkins et al., 2023; Ocasio, 1997). We thus quantify collective attention via time-resolved analyses of conversational content—when does the team talk about what?—using topic timing/synchrony derived from transcripts (Weng et al., 2012; Woolley et al., 2023).
@@
Together with collective attention, shared mental models are the collective cognitive structures that teams develop to interpret and predict their environment (DeChurch and Mesmer-Magnus, 2010). They capture a group’s shared understanding of “what is happening, what is likely to happen next, and why it is happening” (Mohammed et al., 2010, p. 879). Such shared mental models improve motivation, coordination, and performance, particularly when tasks require mutual dependence on diverse skills and expertise (Lewis, 2003). Building on this literature, we operationalize the alignment of shared mental models using a validated and widely used shared mental models survey instrument (Study 2) (Johnson et al., 2007). We also capture alignment of mental models indirectly, drawing on collective framing (Mathieu et al., 2022) so that when subjects utilize team-framed language to describe task execution, they are verbalizing a shared mental model (Study 1).